OpenStudy (alexh107):

Find all solutions to the equation. sin x = square root 3/2

OpenStudy (alexh107):

These were the answer choices

OpenStudy (anonymous):

\[\sin x=\frac{ \sqrt{3} }{ 2 }=\sin(2 n \pi+ \frac{ \pi }{ 3 }),\sin(2 n\pi+ \left( \pi-\frac{ \pi }{ 3 } \right))\] \[x=\left( 2n \pi+\frac{ \pi }{ 3 } \right),\left( 2n \pi+\frac{ 2 \pi }{ 3 } \right)\] where n is an integer.

OpenStudy (alexh107):

Sorry I'm confused, I'm not sure what you did and how that relates to the answer choices.

OpenStudy (anonymous):

third option is same.

OpenStudy (alexh107):

Okay, thank you!

OpenStudy (anonymous):

yw